A water tank in the form of an inverted right-circular cone is being emptied at the rate
of 6m^3 / min. The altitude of the cone is 24m, and the base radius is 12m. Find how fast the water level is lowering when the water is 10m deep?
1
Expert's answer
2020-12-08T17:00:25-0500

Let V be the volume of the water.

dV/dt=6m3/min

Let h be the height of the water

we want dh/dt

Need to find a relationship between dV/dt and dh/dt

V=1/3 πr2h

If V is only the function of h, insted of h and r, we could take derivatives.

By similar triangles r/h= 12/24

r=h/2

"V=\frac{1}{3}π(\frac{h}{2})^2h=\frac{h^3π}{12}"

"\frac{dV}{dt}=\frac{3πh^2}{12}\frac{dh}{dt}=\frac{πh^2}{4}\frac{dh}{dt}"

Where looking for dh/dt when h=10

"6=\frac{π10^2}{4}\frac{dh}{dt}"

"\frac{dh}{dt}=\frac{24}{100π}=\frac{6}{25π}"

Answer:"\frac{6}{25π}" m/min